The work function of potassium metal is 2.30 eV. Monochromatic light is shined on its surface, and photoelectrons with a maximum speed of 7.50E5 m/s are ejected. What is the wavelength of the light?
Answer #1

Inveto = 3.68x10-19) lev o zwork function $ = (2-30ev) bx1only] += (he (6.63x10-34) (3x108 m/s) Ü, 19.1X 10°) (7-5X105)4+ (3